OpenStudy (anonymous):

I will give medal\ A history class is comprised of 10 female and 5 male students. If the instructor of the class randomly chooses 8 students from the class for an oral exam, what is the probability that 5 female students and 3 male students will be selected? Round your answer to 3 decimal places.

OpenStudy (anonymous):

Sure Sweetie

OpenStudy (anonymous):

Use the hypergeometric distribution. In all, there are (14 choose 6) ways of choosing 6 students from the class. Of these, there are (6 choose 2)·(8 choose 4) ways of choosing 2 girls and 4 boys. So the desired probability is ((6 choose 2)·(8 choose 4))/(14 choose 6) ((6!/(4!2!))·(8!/(4!4!))) / (14!/(6!8!)) = (5·10)/(11·13) = 50/143 = 0.3497.

OpenStudy (anonymous):

Did this Help? :)

OpenStudy (anonymous):

Yes